In the aggregate expenditure model for a closed economy, assuming investment, government spending and taxes are exogenous, if the marginal propensity to consume is 0.8, a simultaneous 50 unit increase in government spending and a 20 unit decrease in investment will change equilibrium income by:

Group of answer choices

350 units.

87.5 units.

500 units.

150 units.
1
Expert's answer
2021-09-24T22:18:02-0400

Solution:

The correct answer is 150 units

MPC = 0.8

First, calculate the spending multiplier:

Spending Multiplier = "\frac{1}{1 - MPC} = \frac{1}{1 - 0.8} = \frac{1}{0.2} = 5"


Spending Multiplier = 5

Calculate the change in equilibrium income:

= (50 ΔG "\times" 5) + (20 ΔI "\times" -5) = 250 – 100 = 150 units

The equilibrium income will change by 150 units